# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v1

@@ 4.1 Settings
Simulating hypothesis validation is crucial for our experiments. However, designing individual simulation environments for each experiment is costly. Therefore, we employ surrogate models deployed locally, behaving like a HTTP-based tool for the experiment agent. In our experiments, the task can be “Find a molecule candidate with the highest bio-activity.” To address such tasks, the hypothesis agent proposes hypotheses and the experiment agent validate them using the tool, aiming to identify optimal candidates.
To ensure a fair comparison and focus on core capabilities, all agents utilize QwenMax (Yang2024Qwen2TR, ) as the base LLM and are prohibited from accessing external search tools. The implementation details are in Appendix C.1, and role-playing prompts for all agents are provided in Appendix E. Each experiment was repeated three times with different random seeds, and we report the mean and standard deviation (std) of the relevant metrics for each method.
@@ 4.2 Experimental Scenarios
Nanohelix Optimization (NHO, see Appendix D.1). We developed a high-fidelity surrogate model ($r^{2}=0.98$) to predict g-factor values of a nanohelix material. The model was trained on DFT-simulated data Wu2025MachineLS , enabling efficient exploration of the four-dimensional parameter space (radius of fiber and helix, number of turns, etc.), describing the physical structure of a standard helix.
Molecular Bio-activity Optimization (MBO, see Appendix D.2). We constructed a surrogate model ($r^{2}=0.91$) to predict molecular bio-activity using 50,000 molecule-activity pairs sampled from ChEMBL35 Zdrazil2023TheCD . This model maps SMILES representations directly to pChEMBL values, facilitating rapid in silico screening of candidate therapeutics.
Superconductor Optimization (SPO, see Appendix D.3). We constructed a surrogate model ($r^{2}=0.91$) to predict a superconductor’s $T_{c}$ using 26,321 records of superconductor data publicly available (Hamidieh2018ADS, ). The model learns the map between the compositional features of superconductor to the $T_{c}$ value, facilitating the discovery of room-temperature superconducting materials.
@@ 4.3 Baselines
To evaluate PiFlow’s specific guidance strategy under uncertainty, we therefore benchmark against the following baselines: (1) Reasoning and Acting (ReAct) (Yao2022ReActSR, ). ReAct enables agents to iteratively formulate hypotheses, design/execute experiments, and interpret results, representing a foundational approach to the scientific method loop. (2) Meta Plan Optimization (MPO) (Xiong2025MPOBL, ). MPO utilizes an LLM planner that generates guidance based on past exploration experiences. This offers a contrast to PiFlow’s distinct, explicit mechanism for principled uncertainty reduction. (3) Vanilla Agent System (Vanilla). A Vanilla agent, where a single LLM conducts the discovery process independently, serves as a fundamental benchmark to illustrate the challenges of unguided exploration. While many other advanced agent frameworks exist, their core focus often differs. For instance, Agent-Oriented Planning (AOP) (Li2024AgentOrientedPI, ) is optimized for tasks with known structures rather than initial epistemic uncertainty, and Reason for Future, Act for Now (RAFA) (Liu2023ReasonFF, ) prioritizes predefined task completion over the iterative evidence building from scratch that PiFlow targets. Our evaluation against these baselines is designed to highlight PiFlow’s advantages in discovery efficiency.
@@ 4.4 Evaluation Metrics
We employ two metrics to comprehensively evaluate the performance of PiFlow, as detailed below.
@@
Specifically, for nanomaterial chirality with g-factor value, the theoretical maximum is reported as $\mu_{\text{absolute}}^{\text{g-factor}}=2.0$, indicating a total selectivity for CP light (Greenfield2021PathwaysTI, ), the larger the g-factor, the stronger the chirality; for molecular bio-activity with pChEMBL value, a strict threshold of $\mu_{\text{absolute}}^{\text{pchembl}}=6.5$ has been reported (Lenselink2017BeyondTH, ) for lead compound optimization stage, larger value indicates a better bio-activity and vice versa. for superconductivity Tc, the reference value is set to be $25^{\circ}\mathrm{C}$ , which is equal to 298.15K, denoted as $\mu_{\text{absolute}}^{\text{Tc}}=298.15K$.
Area Under the Curve (AUC). Exploration efficiency requires both (1) rapid convergence and (2) high objective values. We quantify these two factors by defining the AUC metric. Given the trajectory $p_{k},y_{k}\right>\in\mathcal{T}_{t}$ across $t$ steps, AUC is computed via the trapezoidal rule. For meaningful comparisons, we normalize by the maximum possible area:
@@
This normalized metric captures the ability to maintain high performance throughout the optimization trajectory, with higher values indicating more efficient exploration.
@@ 5.1 Performance Comparison
Table 1 demonstrates that PiFlow achieves significant improvements over all baselines across three benchmarks of NHO, MBO, and SPO. In terms of achieving the target property, measured by SQ, PiFlow consistently leads, outperforming ReAct, MPO, and Vanilla systems by an average of approximately 207.6%, 34.1%, and 94.1%. Furthermore, PiFlow demonstrates superior exploration efficiency, evidenced by its leading AUC scores, translating to average AUC improvements of approximately 146.5% over ReAct, 73.6% over Vanilla and 54.0% over MPO. Additionally, we compare the exploration trajectory regarding the experiment outcome, as presented in Figure 3. It shows that, PiFlow is always the fastest one in reaching the best solution while exploring. Although there are some fluctuations in the process, the general upward trend continues to surpass other methods.
These significant advancements in both SQ and AUC strongly support our central claim: the strategic and principled guidance provided by the PiFlow component, rooted in its principle-aware optimization, effectively steers the multi-agent hypothesis-validation loop towards more rapid and fruitful scientific discovery.
@@ 5.2 Ablation Study
We conduct several ablations to evaluate the components of PiFlow. The studies investigate the impact of PiFlow itself, the choice of foundation LLMs, the influence of a key hyperparameter $\lambda$, which is for balancing exploitation and exploration, and the effect of the Thought mode in LLM reasoning. For these studies, conducted on the NHO task, performance is evaluated based on the same metrics, AUC (%) and SQ (%).
Impact of PiFlow Integration (Plug-and-Play). To isolate the direct benefit of PiFlow, we compared the performance of two different LLMs, GPT4.1-mini and Qwen3-32B, w/ and w/o the PiFlow component. This is achieved by only including or excluding the steered principle with one of the three specified action types to the Planner Agent via prompt, which then directs subsequent hypothesizing and validation.
@@
As shown in Table 2, for GPT4.1-mini, integrating PiFlow increases the AUC from 37.12% to 41.68% and substantially boosted the SQ from 40.14% to 66.38%. This represents an approximate 12.3% improvement in AUC and a significant 65.4% improvement in SQ. Similarly, for the Qwen3-32B model, the inclusion of PiFlow improves AUC from 27.04% to 37.51% (a 38.7% increase) and SQ from 54.84% to 58.76% (a 7.1% increase). $\diamond$ Takeaway: These findings underscore the significant positive impact of PiFlow on both AUC and SQ, confirming its effectiveness as a Plug-and-Play enhancement for MAS in scientific discovery.
Influence of Foundation Models. The choice of underlying LLMs may significantly affect the reasoning process, thereby affecting the performance. To evaluate the model influence while balancing the agent functions, we replace LLMs of $\mathcal{A}_{H}$ and $\mathcal{A}_{P}$ evaluation. The $\mathcal{A}_{E}$responsible for tool usage, consistently used QwenMax to ensure functional tool interaction. We evaluate several state-of-the-art LLMs (e.g., Claude-3.7-sonnet anthropic2025 , GPT4.1-mini openai2025 , Gemini-2.5-pro-exp-0325 google2025 , Qwen3-32B and QwenMax yang2024qwen2 ) on the NHO task, and the results are summarized in Table 3.
Table 3: Model Performance Comparison
Method/Setting AUC (%) SQ (%)
Claude-3.7-sonnet 38.60 $\pm$4.30 78.50 $\pm$3.74
GPT4.1-mini 41.68 $\pm$17.91 66.38 $\pm$14.90
Gemini-2.5-pro-exp-03-25 28.43 $\pm$13.81 69.64 $\pm$17.10
Qwen3-32B 37.51 $\pm$7.70 58.76 $\pm$6.18
QwenMax 63.51 $\pm$11.18 76.82 $\pm$4.54
Among these models, QwenMax demonstrates the highest AUC at 63.51% and a strong SQ of 76.82%. Claude-3.7-sonnet achieves the highest SQ at 78.50% with an AUC of 38.60%. Other models like GPT4.1-mini (AUC 41.68%, SQ 66.38%) and Qwen3-32B (AUC 37.51%, SQ 58.76%) also show competent performance. Gemini-2.5-pro-exp-03-25 yields an AUC of 28.43% and an SQ of 69.64%. $\diamond$ Takeaway: The performance variation across different models highlights the importance of model selection for specific tasks within the PiFlow-enhanced MAS.
Sensitivity to Hyperparameter $\lambda$. We investigate the impact of $\lambda$ in Eq. 1. The specific role of $\lambda$ is to balance exploration and exploitation, with larger $lambda$ value places greater emphasis on exploration. The results conducted on the NHO task using QwenMax model for different values of $\lambda$ are detailed in Table 4.
Table 4: Lambda Ablations
Setting AUC (%) SQ (%)
$\lambda=0.1$ 41.32 $\pm$5.90 66.45 $\pm$9.49
$\lambda=0.3$ 44.28 $\pm$2.83 66.43 $\pm$9.28
$\lambda=0.5$ 32.99 $\pm$11.16 59.02 $\pm$3.56
$\lambda=0.7$ 40.50 $\pm$2.89 56.40 $\pm$4.79
$\lambda=0.9$ 34.57 $\pm$10.33 62.49 $\pm$8.38
From the table, AUC varied with different $\lambda$ values, peaking at 44.28% when $\lambda=0.3$. The SQ remained relatively high for $\lambda=0.1$ (66.45%) and $\lambda=0.3$ (66.43%), but showed more variability with other settings. For instance, $\lambda=0.5$ resulted in a lower AUC (32.99%) and SQ (59.02%). While increasing $\lambda$, AUC tends to decrease, as a stronger emphasis on exploration can lead to more varied hypothesis selection. These results suggest that the system’s performance, particularly its exploration efficiency, is sensitive to the choice of $\lambda$, with $\lambda=0.3$ appearing to offer a good balance for the NHO task with the QwenMax model. $\diamond$ Takeaway: The results suggest that $\lambda$ indeed effects the performance in both AUC and SQ, fine-tuning $\lambda$ can be crucial for optimizing performance.
Effect of Thought Mode. We also conducted an ablation study on an internal thought mode of the LLM (referred to as Think in Table 5) on the NHO benchmark. This is for agents based on Qwen3-32B and Qwen3-8B models, which support ON/OFF <think> token generation with system prompt including or excluding /no_think. This mode is intended to enable more explicit reasoning steps.
@@
Interestingly, for both Qwen3-32B and Qwen3-8B, disabling the Thought Mode led to improved performance. For Qwen3-32B, removing the Thought Mode increased AUC from 37.51% to 45.51% and SQ from 58.76% to 68.86%. Similarly, for Qwen3-8B, performance without the Thought Mode was better, with AUC rising from 30.55% to 42.09% and SQ increasing from 54.49% to 61.59%. $\diamond$ Takeaway: These counter-intuitive results suggest that the current implementation or utility of this specific Thought Mode may introduce overhead or misguide the agents in these contexts, warranting further investigation into its design and application.
@@ 5.3 Theoretical Alignment
To empirically validate PiFlow’s theoretical guarantees (Theorem 1), we analyze key aspects of its exploration dynamics, including (1) the bound of average regret in PiFlow and (2) the relationship between regret and information gain, as shown in Figure 4.
(Theoretical Prediction 1) Average regret decay with $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$. Figure 4a presents the average regret as a function of iterations on a log-log scale. The empirical results demonstrate a strong concordance with the theoretical prediction of $\mathcal{O}(\sqrt{T})$ sublinear cumulative regret. This is evidenced by average regret trajectories adhering to the $T^{-0.5}$ decay (fitted $c\cdot T^{-0.5}$ with $c=1.37,R^{2}=0.96$), a pattern most runs consistently follow. Notably, one run (ID=1), after an initial sharp regret decrease, shows a transient increase before resuming decay. This illustrates PiFlow’s robust exploration avoiding potential local optima with a characteristic of its Min-Max strategy.
(Theoretical Prediction 2) As information gain decreases, the expected regret also decreases. Theorem 1 also posits that decreasing information gain correlates with decreasing expected regret. Figure 4b empirically verifies this with PiFlow. The scatter plot of regret versus information gain (points colored by iteration) reveals a clear positive association, confirmed by a fitted trend line: lower information gain generally corresponds to lower regret. Early iterations (higher information gain and regret) transition to later iterations (lower information gain and regret). This empirical trend strongly supports the theoretical analysis, showcasing PiFlow’s effective shift from exploration to exploitation.
@@
(c) NHO Exploration Trajectory
Figure 4: Empirical Validation of PiFlow’s Theoretical Alignment. (a) Average regret over iterations aligns with the theoretical sublinear decay; (b) Scatter plot with positive trend of observed regret versus information gain, supporting theoretical expectations; (c) An exemplary PiFlow exploration trajectory.
Empirical Validation of Exploration Strategy on NHO Landscape. To further illustrate PiFlow’s operational dynamics, Figure 4c visualizes an exemplary exploration trajectory on the NHO task. The objective is to identify nanohelix parameters yielding the maximum property value. The background displays the NHO objective landscape by Principal Component Analysis (PCA), where contours indicate g-factor values fitted over the source data.
Aligning with its robust design, PiFlow’s trajectory shows initial broad exploration (iterations 1-16), then successful navigation of a low-quality “valley” (iterations 16-21) crucial for avoiding local optima, before efficiently converging to a high-g-factor region and identifying near-optimal parameters (iterations 21-24).
$\diamond$ Takeaway: The above regret analysis and visualized behavior empirically support the claim that PiFlow’s principle-aware mechanism translates its theoretical guarantees into effective and efficient scientific discovery.
6 Limitation
While PiFlow shows notable improvements through its principled Min-Max optimization, its practical implementation approximates a key theoretical component. This means the current system may not fully capture all nuances of true, model-based information gain, especially the direct adversarial interplay with all possible manifestations of the unknown evaluation function $f^{*}$ from the theoretical objective. Future research could explore more direct estimations of mutual information for this heuristic within the PiFlow framework to potentially further enhance its strategic guidance.

### v2

@@ 4.1 Settings
To rigorously evaluate the effectiveness and versatility of our PiFlow framework, we conducted experiments across three distinct scientific discovery scenarios. While direct hypothesis validation in real-world labs is prohibitively expensive, we employ high-fidelity surrogate models deployed locally, which serve as the primary evaluation tool. Across all scenarios (Section 4.2), we frame the scientific discovery challenge as a unified task: “Find a candidate in a complex parameter space that maximizes a target property (e.g., bio-activity).
To ensure a fair comparison and focus on core capabilities, all agents utilize QwenMax (Yang et al., 2024a) as the base LLM and are prohibited from accessing external search tools. The complete experimental setup is detailed in Appendix S, and the role-playing prompts for PiFlow are provided in Appendix T.
@@ 4.2 Experimental Scenarios
To comprehensively assess PiFlow’s performance, we design three scenarios that represent canonical challenges in scientific exploration: optimization in continuous, discrete, and mixed parameter spaces:
Nanohelix Optimization (NHO). We use a surrogate model ($r^{2}=0.98$) trained on DFT-simulated data following Wu et al. (2025) to predict nanohelix chirality from four continuous geometric parameters, enabling efficient exploration of the design space (Appendix S.1).
Molecular Bio-activity Optimization (MBO). We build a surrogate model ($r^{2}=0.91$) to predict bio-activity from SMILES strings, trained on 50,000 molecules from ChEMBL35 (Zdrazil et al., 2023). This facilitates high-throughput screening in a discrete chemical space. (Appendix S.2).
Superconductor Optimization (SPO). Following  Hamidieh (2018), we train a surrogate model ($r^{2}=0.91$) to map a material’s mixed continuous and discrete compositional features to its critical temperature ($T_{c}$), accelerating the discovery of room-temperature superconductors. (Appendix S.3).
@@ 4.3 Baselines
To evaluate the strategic guidance of PiFlow under uncertainty, we therefore benchmark against the following baselines: (1) Reasoning and Acting (ReAct) (Yao et al., 2022). ReAct enables agents to iteratively formulate hypotheses, design/execute experiments, and interpret results. It represents a foundational approach to structured reasoning. (2) Meta Plan Optimization (MPO) (Xiong et al., 2025). MPO employs a trained LLM planner that provides high-level, general guidance. This serves as a direct counterpoint to PiFlow’s explicit, structured mechanism for principled uncertainty reduction. (3) Vanilla Agent System (Vanilla). This baseline consists of an MAS operating without any principled strategic oversight. It is intended to establish a performance floor, demonstrating the limitations of unguided exploration reliant solely on agent role-playing.
While other advanced frameworks exist, such as Agent-Oriented Planning (AOP) (Li et al., 2024) and Reason for Future, Act for Now (RAFA) (Liu et al., 2023), their design objectives diverge from the scope of our evaluation. AOP is optimized for tasks with well-defined structures rather than initial epistemic uncertainty, while RAFA focuses on completing predefined goals rather than the iterative, de novo evidence-gathering process central to PiFlow. Therefore, our curated selection of baselines is specifically designed to isolate and evaluate our contribution to discovery efficiency in uncertain environments.
@@ 4.4 Evaluation Metrics
Two metrics are employed to evaluate the performance of PiFlow, as detailed below:
@@
Specifically, for NHO, the theoretical maximum is reported as $\mu_{\text{absolute}}^{\text{g-factor}}=2.0$ following Greenfield et al. (2021). The larger the g-factor, the stronger the chirality. For MBO, a strict threshold of $\mu_{\text{absolute}}^{\text{pchembl}}=6.5$ has been reported (Lenselink et al., 2017) for lead compound optimization stage, larger value indicates a higher bio-activity and vice versa. For SPO, the reference value is set to be $25^{\circ}\mathrm{C}$ , which is equal to room temperature 298.15K, denoted as $\mu_{\text{absolute}}^{\text{Tc}}=298.15K$.
Area Under the Curve (AUC). Exploration efficiency requires both (1) rapid convergence and (2) high objective values. We quantify these two factors by defining the AUC metric. Given the trajectory $\left<p_{k},y_{k}\right>\in\mathcal{T}_{t}$ across $t$ steps, AUC is computed via the trapezoidal rule. For meaningful comparisons, we normalize by the maximum possible area:
@@
In summary, SQ measures the quality of the final outcome, while AUC evaluates the entire discovery process by rewarding both speed and consistency. The detailed rationale for these metric designs is provided in Appendix R.2.
@@ 5.1 Performance Comparison
Table 1 demonstrates that PiFlow achieves significant improvements over all baselines across three benchmarks of NHO, MBO, and SPO. In terms of achieving the target property, measured by SQ, PiFlow consistently leads, outperforming ReAct, MPO, and Vanilla systems by an average of approximately 207.6%, 34.1%, and 94.1%. Additionally, as shown in Figure 3, PiFlow is always the fastest one in reaching the best solution while exploring. More details about models’ response analysis can be seen at Appendix I.
It is worth noting the considerable std observed in our results. This variance is inherent to the stochastic nature of LLM-based agents and can be attributed to the phenomenon of cumulative error. An early suboptimal decision or flawed hypothesis can set an agent on an inefficient discovery trajectory, with reasoning errors compounding over subsequent iterations. This instability highlights why relying solely on the final SQ can be insufficient. In contrast, the AUC provides a more holistic performance measure. A high AUC, as consistently demonstrated by PiFlow, indicates that the agent not only found a high-quality solution but also maintained a robust and efficient path, thereby successfully mitigating the primary risk of cumulative error.
Takeaway: PiFlow substantially outperforms all baselines in SQ and exploration efficiency (AUC). Its strategic robustness is evidenced by its high AUC, which mitigates the cumulative errors of LLMs, and by its success from modest starting points that simulate realistic initial uncertainty (see Appendix P.2).
@@ 5.2 Ablation Study
We conduct several ablations to evaluate PiFlow. For these studies, conducted on the NHO task, performance is evaluated based on the same metrics, AUC (%) and SQ (%).
Plug-and-Play. To isolate the direct benefit of PiFlow, we compared the performance with two different LLMs, GPT4.1-mini and Qwen3-32B under the setting of w/ and w/o PiFlow. This is achieved by only including or excluding the steered principle to the Planner Agent via prompt, which then directs subsequent hypothesizing and validation.
As shown in Table 2, for GPT4.1-mini, integrating PiFlow increases the AUC from 37.12% to 41.68% and substantially boosted the SQ from 40.14% to 66.38%. This represents an approximate 12.3% improvement in AUC and a significant 65.4% improvement in SQ. Similarly, for the Qwen3-32B model, the inclusion of PiFlow improves AUC from 27.04% to 37.51% (a 38.7% increase) and SQ from 54.84% to 58.76% (a 7.1% increase).
@@
Thought Mode Effect. We also conduct ablations on the internal thought mode of the LLM (referred to as Think in Table 3). This is for agents based on Qwen3-32B and Qwen3-8B models, which support ON/OFF <think>...</think> generation with system prompt including or excluding /no_think. This mode is intended to enable more explicit reasoning steps.
@@
Interestingly, for both Qwen3-32B and Qwen3-8B, disabling the Thought Mode leads to improved performance, as shown at Table 3. We hypothesize that this phenomenon is due to cognitive fixation. As the key issue is, how do LLM propose scientific hypothesis, forcing the LLM to generate an explicit Chain-of-Thought may cause it to commit prematurely to its own initial line of reasoning. If the first step in its logic is flawed, the entire chain can be led astray, creating a cognitive fixation that is hard to escape. In contrast, disabling the think mode may force the model to rely more on its powerful, holistic pattern-matching capabilities, allowing it to make more intuitive leaps directly from the data (PiFlow’s guidance and the experimental history) to a hypothesis, bypassing potentially flawed intermediate reasoning steps.
Takeaway: Our ablations confirm PiFlow is a robust, plug-and-play enhancement that consistently boosts performance across models. Concurrently, we find that an agent’s internal reasoning is critical: disabling explicit "Think" modes paradoxically improves performance, suggesting that avoiding cognitive fixation allows the LLM to better leverage holistic pattern-matching. This highlights the synergy between high-level strategic guidance from PiFlow and internal reasoning.
5.3 Further Analyses of Robustness
We conducted extensive analyses to probe the robustness and practical utility of PiFlow. The key findings are as follows:
Recovery from poor initialization. The system demonstrates resilience by successfully recovering from deliberately incorrect initial principles, eventually matching the performance of a run guided by correct initialization. This underscores the robustness of the Min-Max (Appendix P).
Temporal dynamics of principle evaluation. PiFlow continuously re-evaluates the utility of scientific principles as new evidence is gathered. This allows it to dynamically discard initially promising but ultimately flawed avenues while elevating principles that prove more fruitful later, showcasing an adaptive balance between exploration and exploitation (Appendix N).
Superior performance against numerical search. On complex tasks with ill-defined search spaces, PiFlow substantially outperforms Bayesian Optimization, achieving an SQ of 84.55% versus BO’s 38.15% in the MBO task, without requiring manual parameter space engineering (Appendix J).
Seamless plug-and-play integration. We validate the practicality of PiFlow through a successful integration with ChemToolAgent (Yu et al., 2024), guiding it to a high-value molecule (pChEMBL of 5.90) without requiring any architectural modifications (Appendix K).
Manageable computational complexity. The core decision mechanism of PiFlow has a computational complexity of $\mathcal{O}(t^{2}\cdot d)$, where $t=|\mathcal{T}_{t}|$ and $d$ is the embedding dimension. This could be optimized to near-linear time for large-scale tasks (Appendix L).
Cost-effectiveness. Superior performance is achieved cost-effectively, reducing token consumption by up to 27% compared to the Vanilla Agent. PiFlow constitutes only 1.5% tokens (Appendix M).
Generalizability and controllability. PiFlow is compatible with various LLMs, and its behavior can be tuned via the $\lambda$ hyperparameter, for which we provide a clear heuristic (Appendices O, Q).
@@ 5.4 Theoretical Alignment
To empirically validate the theoretical guarantees (Theorem 1) of PiFlow, we analyze key aspects of its exploration dynamics, including (1) the bound of average regret and (2) the relationship between regret and information gain, as shown in Figure 4.
(Theoretical Prediction 1) Average regret decay with $\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)$. Figure 4a presents the average regret as a function of iterations on a log-log scale. The alignment evidenced by average regret trajectories adhering to the $T^{-0.5}$ decay (fitted $c\cdot T^{-0.5}$ with $c=1.37,r^{2}=0.96$), a pattern most runs consistently follow. Notably, one run (ID=1), after an initial sharp regret decrease, shows a transient increase before resuming decay. This illustrates PiFlow’s robust exploration avoiding potential local optima with a characteristic of its Min-Max strategy.
(Theoretical Prediction 2) As information gain decreases, the expected regret also decreases. As shown in Figure 4b, the scatter plot of regret versus information gain (points colored by iteration) reveals a clear positive association, confirmed by a fitted trend line: lower information gain generally corresponds to lower regret. Early iterations (higher information gain and regret) transition to later iterations (lower information gain and regret).
@@
(c) Exploration Trajectory
Figure 4: Empirical Validation of PiFlow’s Theoretical Alignment. (a) Average regret over iterations aligns with the theoretical sublinear decay; (b) Scatter plot with positive trend of observed regret versus information gain, supporting theoretical expectations; (c) An exemplary PiFlow exploration trajectory in NHO task.
Empirical validation of exploration strategy. Figure 4c displays the trajectory of PiFlow on the NHO objective landscape, which is visualized via Principal Component Analysis (PCA) with contours indicating g-factor values. The path demonstrates a principled strategy: initial broad exploration (iterations 1-16), followed by navigating a low-quality "valley" to escape a local optimum (iterations 16-21), and culminating in efficient convergence to a high-g-factor region (iterations 21-24). This progression empirically validates how our theoretical design translates into effective discovery.
Takeaway: Our empirical analysis corroborates the theoretical guarantees of PiFlow, demonstrating that its strategy is both principled and effective. The alignment of regret dynamics with the predicted sublinear rate and its coupling with information gain validates a robust mechanism that translates theoretical efficiency into a practical ability to escape local optima and converge reliably.